OpenStudy (lakers):

Solve the quadratic equation. 4x2 - 30 = 0 A. ± √15 B. ± √15/√2 C. ± √15 /2 D. ± √30 /2

OpenStudy (anonymous):

quadratic formula \[\frac{ -b \pm \sqrt{b^2-4*a*c} }{ 2a }\]

OpenStudy (anonymous):

ax^2+bx+c=0 is where that expression comes from (if that makes sense)